Grade 9Engineering Entrance Exams

Difference between mirror reflected light rays and rays of a screen

rishav kumar
12 Years agoGrade 9

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

Distance to the image is the big difference.

The reflected image in a plane mirror is composed of rays that appear to come from a pointbehindthe mirror (by the same distance that the image that is being reflected is in front of the mirror:d0=did0=di). For non-plane mirrors you'll have to work a little harder, but locally spherical bits can be treated with the usual1/f=1/di+1/do.1/f=1/di+1/do.The light rays emanating from a screen come from the surface of the screen.

Regards

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago